System of Linear Equation Word Problem : the sum of two numbers 111 and their difference is 45.
Answer by rothauserc(4718)

I make the initial condition x > y
:
x + y = 111
x - y = 45
:
solve second equation for x
:
x = y + 45
:
substitute for x in first equation
:
y+45 +y = 111
2y = 66
:
******************
y = 33 and x = 78
******************